OpenStudy (anonymous):

Simplify and write the answer in standard form 4+3i/ 5+2i

zepdrix (zepdrix):

Hi there, \(\Large\bf \color{#008353}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

We want to multiply the top and bottom by the conjugate of our denominator, \[\Large\bf\sf \frac{4+3\mathcal i}{5+2\mathcal i}\color{royalblue}{\left(\frac{5-2\mathcal i}{5-2\mathcal i}\right)}\]

zepdrix (zepdrix):

Remember your rule for multiplying complex conjugates?\[\Large\bf\sf (a-b\mathcal i)(a+b\mathcal i)\quad=\quad a^2+b^2\]

zepdrix (zepdrix):

So we get,\[\Large\bf\sf =\quad \frac{(4+3\mathcal i)(5-2\mathcal i)}{5^2+2^2}\]

zepdrix (zepdrix):

Do you understand how to proceed from there?

OpenStudy (anonymous):

can you continue please

zepdrix (zepdrix):

Nooo you gotta do some of the work here silly!! :) Do you remember how to multiply brackets out? Does FOIL sound familiar maybe?

OpenStudy (anonymous):

i know how to foil so my top part will look like (4+3i)(4-3i)(5+2i)(5-2i)

zepdrix (zepdrix):

No, your top is,\[\Large\bf\sf (4+3\mathcal i)(5-2\mathcal i)\]And you need to FOIL that out.

OpenStudy (anonymous):

but i thought every complex number came with its other pair(+or-)

zepdrix (zepdrix):

No.. hmm not sure where you're getting that from :D All complex `roots` come in pairs, yes. So when you're solving an equation for x or something, then yes you'll always get the conjugate along with your complex root. But not when we're just straight up dealing with complex numbers.

OpenStudy (anonymous):

ooooohh sorry i see now so my top answer is -6i^2+7i+20

OpenStudy (anonymous):

all over 29

zepdrix (zepdrix):

Mmmm ok looks good so far!

zepdrix (zepdrix):

Simplify the i^2 term. You should be able to combine it with something else after that.

OpenStudy (anonymous):

so its 7i+26

zepdrix (zepdrix):

Ok looks good so far :) \[\Large\bf\sf \frac{26+7\mathcal i}{29}\]

zepdrix (zepdrix):

Now we want to get it into standard form.

OpenStudy (anonymous):

i thought tha was it?

zepdrix (zepdrix):

We want to split this into two fractions, the real and imaginary parts.

OpenStudy (anonymous):

"but i thought every complex number came with its other pair(+or-)" this is correct when rationalizing Ex. 2i/2i-5 2i(2i+5)/(2i-5)(2i+5)

zepdrix (zepdrix):

I don't know what this is called, but I like to call it the jelly bean method :)